# UniMark: Artificial Intelligence Generated Content Identification Toolkit

### v2

@@ Abstract
The rapid proliferation of Artificial Intelligence Generated Content has precipitated a crisis of trust and urgent regulatory demands. However, existing identification tools suffer from fragmentation and a lack of support for visible compliance marking. To address these gaps, we introduce the UniMark, an open-source, unified framework for multimodal content governance. Our system features a modular unified engine that abstracts complexities across text, image, audio, and video modalities. Crucially, we propose a novel dual-operation strategy, natively supporting both Hidden Watermarking for copyright protection and Visible Marking for regulatory compliance. Furthermore, we establish a standardized evaluation framework with three specialized benchmarks (Image/Video/Audio-Bench) to ensure rigorous performance assessment. This toolkit bridges the gap between advanced algorithms and engineering implementation, fostering a more transparent and secure digital ecosystem. The code is available at https://github.com/AI-LAB-C-for-S-T-AI/AIGC-Identification-Toolkit.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) and generative diffusion models have made remarkable progress. From text-domain models such as GPT-5 [10] and Qwen [1], to video generation models like Sora 2 [13], Wan 2.1 [16], and Nano Banana [8], Artificial Intelligence Generated Content (AIGC) is reshaping the production of digital content at an unprecedented pace. However, the widespread adoption of these technologies has also precipitated a profound crisis of trust. The abuse of Deepfake technology, the viral spread of fake news, and increasingly severe copyright infringement issues are eroding the foundation of public trust in digital information. Facing this challenge, establishing a reliable mechanism for Content Identification and Tracing is no longer optional but an urgent necessity for the digital ecosystem. Governments worldwide have also attached great importance to this issue; for instance, China’s national standard Information Security Technology – Method for Identifying Content Generated by Artificial Intelligence [15] and the European Union’s AI Act [4] have both set forth clear specifications and requirements for the identification of AIGC.
Existing AIGC identification solutions fall primarily into two categories: In-processing (generation process embedding) and Post-processing (generation after-treatment). In-processing methods embed watermarks directly during the model inference stage. Representative works include SynthID [3] and GumbelSoft [7] in the text domain; Stable Signature [5], Tree-Ring [17], and PRC [9] in the image domain; and Groot [12] in the audio domain. Although these methods theoretically offer high imperceptibility, they face significant limitations in practical application: they typically require access to internal model parameters (such as logits), making them inapplicable to black-box APIs like GPT-5 or Midjourney, and they cannot handle pre-existing generated data. Consequently, to build a universal and model-agnostic solution, Post-processing schemes have become a more practical choice. Related research, such as PostMark [2] for text, ZoDiac [18] for images, VideoSeal [6] for video, and AudioSeal [14] for audio, has demonstrated immense potential. However, the current ecosystem of identification tools suffers from severe fragmentation—State-of-the-Art (SOTA) algorithms are scattered across different codebases with inconsistent interface standards, making integration difficult. More importantly, existing tools mostly focus solely on implicit technical watermarks, neglecting “Visible Marking”—a critical dimension for regulatory compliance—thus failing to meet the requirements for user right-to-know mandated by regulations such as the EU AI Act.
@@
Finally, considering the rapid iteration of AIGC algorithms, we designed the architecture as a high-extensibility algorithm platform. By defining standardized interface protocols, researchers can integrate the latest SOTA algorithms, such as PostMark [2] and CredID [11], into the framework at zero cost, enabling it to evolve continuously with technological developments.
@@ 2.1 Unified Engine Architecture
  • Visible Marking (operation=’visible_mark’): Applies explicit compliance markers (e.g., visual overlays or audio prompts) to meet regulatory transparency requirements, such as those mandated by the EU AI Act [4].
@@ 2.2 Multimodal Algorithm Integration
For both image and video modalities, we unify the backend using VideoSeal [6]. By treating images as single-frame videos, we leverage a single pre-trained encoder to embed watermarks that are robust against both spatial editing (e.g., cropping, resizing) and temporal manipulation (e.g., frame dropping). This approach significantly reduces code redundancy and maintenance overhead while ensuring consistent performance across visual content. Furthermore, we construct an automated pipeline that seamlessly links generative models with the watermarking process; for instance, video content generated by Wan 2.1 [16] is automatically passed to the VideoSeal encoder for immediate protection.
Regarding other modalities, we integrate AudioSeal [14] for audio watermarking, which enables precise identification of AI-generated segments within longer streams through localized detection. For text, we employ PostMark [2] to address the prevalence of closed-source LLMs. Unlike generation-time methods that require access to model logits, PostMark operates as a post-processing step, ensuring our toolkit remains model-agnostic and compatible with black-box APIs such as GPT-4.

### v3

@@ Abstract
The rapid proliferation of Artificial Intelligence Generated Content has precipitated a crisis of trust and urgent regulatory demands. However, existing identification tools suffer from fragmentation and a lack of support for visible compliance marking. To address these gaps, we introduce the UniMark, an open-source, unified framework for multimodal content governance. Our system features a modular unified engine that abstracts complexities across text, image, audio, and video modalities. Crucially, we propose a novel dual-operation strategy, natively supporting both Hidden Watermarking for copyright protection and Visible Marking for regulatory compliance. Furthermore, we establish a standardized evaluation framework with three specialized benchmarks (Image/Video/Audio-Bench) to ensure rigorous performance assessment. This toolkit bridges the gap between advanced algorithms and engineering implementation, fostering a more transparent and secure digital ecosystem. The code is available at https://github.com/AI45Lab/UniMark.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) and generative diffusion models have made remarkable progress. From text-domain models such as GPT-5 [gpt4o] and Qwen [qwen2], to video generation models like Sora 2 [sora], Wan 2.1 [wan], and Nano Banana [nano], Artificial Intelligence Generated Content (AIGC) is reshaping the production of digital content at an unprecedented pace. However, the widespread adoption of these technologies has also precipitated a profound crisis of trust. The abuse of Deepfake technology, the viral spread of fake news, and increasingly severe copyright infringement issues are eroding the foundation of public trust in digital information. Facing this challenge, establishing a reliable mechanism for Content Identification and Tracing is no longer optional but an urgent necessity for the digital ecosystem. Governments worldwide have also attached great importance to this issue; for instance, China’s national standard Information Security Technology – Method for Identifying Content Generated by Artificial Intelligence [gb] and the European Union’s AI Act [eu_ai] have both set forth clear specifications and requirements for the identification of AIGC.
Existing AIGC identification solutions fall primarily into two categories: In-processing (generation process embedding) and Post-processing (generation after-treatment). In-processing methods embed watermarks directly during the model inference stage. Representative works include SynthID [synthid] and GumbelSoft [gumbelsoft] in the text domain; Stable Signature [stable_signature], Tree-Ring [tree_rings], and PRC [prc] in the image domain; and Groot [groot] in the audio domain. Although these methods theoretically offer high imperceptibility, they face significant limitations in practical application: they typically require access to internal model parameters (such as logits), making them inapplicable to black-box APIs like GPT-5 or Midjourney, and they cannot handle pre-existing generated data. Consequently, to build a universal and model-agnostic solution, Post-processing schemes have become a more practical choice. Related research, such as PostMark [postmark] for text, ZoDiac [zodiac] for images, VideoSeal [videoseal] for video, and AudioSeal [audioseal] for audio, has demonstrated immense potential. However, the current ecosystem of identification tools suffers from severe fragmentation—State-of-the-Art (SOTA) algorithms are scattered across different codebases with inconsistent interface standards, making integration difficult. More importantly, existing tools mostly focus solely on implicit technical watermarks, neglecting “Visible Marking”—a critical dimension for regulatory compliance—thus failing to meet the requirements for user right-to-know mandated by regulations such as the EU AI Act.
@@
Finally, considering the rapid iteration of AIGC algorithms, we designed the architecture as a high-extensibility algorithm platform. By defining standardized interface protocols, researchers can integrate the latest SOTA algorithms, such as PostMark [postmark] and CredID [credid], into the framework at zero cost, enabling it to evolve continuously with technological developments.
@@ 2.1 Unified Engine Architecture
  • Visible Marking (operation=’visible_mark’): Applies explicit compliance markers (e.g., visual overlays or audio prompts) to meet regulatory transparency requirements, such as those mandated by the EU AI Act [eu_ai].
@@ 2.2 Multimodal Algorithm Integration
For both image and video modalities, we unify the backend using VideoSeal [videoseal]. By treating images as single-frame videos, we leverage a single pre-trained encoder to embed watermarks that are robust against both spatial editing (e.g., cropping, resizing) and temporal manipulation (e.g., frame dropping). This approach significantly reduces code redundancy and maintenance overhead while ensuring consistent performance across visual content. Furthermore, we construct an automated pipeline that seamlessly links generative models with the watermarking process; for instance, video content generated by Wan 2.1 [wan] is automatically passed to the VideoSeal encoder for immediate protection.
Regarding other modalities, we integrate AudioSeal [audioseal] for audio watermarking, which enables precise identification of AI-generated segments within longer streams through localized detection. For text, we employ PostMark [postmark] to address the prevalence of closed-source LLMs. Unlike generation-time methods that require access to model logits, PostMark operates as a post-processing step, ensuring our toolkit remains model-agnostic and compatible with black-box APIs such as GPT-4.